OpenStudy (hockeychick23):

Probability theory predicts that there is 77.6% chance of a particular soccer player making 2 penalty shots in a row. If the soccer player taking 2 penalty shots is simulated 2500 times, in about how many simulations would you expect at least 1 missed shot?

OpenStudy (perl):

do you need help with this

OpenStudy (perl):

be right back

OpenStudy (hockeychick23):

ok and yep i do thanks

OpenStudy (perl):

Probability theory predicts that there is 77.6% chance of a particular soccer player making 2 penalty shots in a row. If the soccer player taking 2 penalty shots is simulated 2500 times, in about how many simulations would you expect at least 1 missed shot? First off , lets get probability. we can use the complement P ( at least 1 missed shot ) = 1 - P ( no missed shots)

OpenStudy (hockeychick23):

ok so its p^2= .776 and p=.88090862182

OpenStudy (hockeychick23):

oh i did something wrong then sorry

OpenStudy (perl):

P ( at least 1 missed shot ) = 1 - P ( no missed shots) = 1 - .776

OpenStudy (perl):

i think you solved for p, which is another way to do it

OpenStudy (perl):

but it might be harder approach

OpenStudy (hockeychick23):

oh ok then, i got p=.224 now

OpenStudy (perl):

right

OpenStudy (perl):

now you can multiply that probability by the total population

OpenStudy (hockeychick23):

ok .224(2500)= 560

OpenStudy (perl):

correct

OpenStudy (hockeychick23):

oh wow that was actually a lot quicker of a problem than i thought, thanks so much!